Biosphere 2 Research Wiki » xe.blog.code.title » 2016126100540 East +4 flow meter swap

2016126100540 East +4 flow meter swap

Last modified by Michael Sibayan on 2016/12/06 10:11
Comments (0) · Attachments (0) · History

Dec 06 2016

On Dec. 5, 2016 from 13:45-13:55, Nate changed out the PE102-75 flow meter with a PE102-38. Looking at the raw counter data it seems to be working. This was a used ebay purchase. 

Data before this time used incorrect scaling factor (1000, instead of the correct 500 for a -75).

Scaling factor was not changed at this point, since the correct value is already in place.

Tags:
Created by Michael Sibayan on 2016/12/06 10:11

This wiki is licensed under a Creative Commons 2.0 license
XWiki Enterprise 6.0.1 - Documentation